An elevator mass of 1700 kg falls from a height of 40 m after a sudden failure in the hoisting cable. The mass is stopped by a spring at the bottom of the shaft having a spring constant of 75 kN/m and spring's compression of 4.44589 m. Determine the acceleration in m/s2 that the elevator undergoes when brought to rest by the spring.
Answer #1

m*g*h = 1/2*m*v^2

v = sqrt(2*g*h)

= sqrt(2*9.8*40)

= 28 m/s

If it is stopped in 4.44589 m distance

v^2 = 2*a*s

28^2 = 2*a*4.44589

a = 88.17 m/s^2 ---> Answer